The dissociation energy of H2 is 430.53 KJ/mol. If H2 is exposed to radiant energy of wavelength 253.7 nm, what % of radiant energy will be converted into K.E.


Grade:

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

dissociation energy is 430.53 KJ/mol

photoon energy=hc/wavelength

Now calculate the radiant energy.....

vikas askiitian expert
13 years ago

bond dissociation energy = 430.53kj/mol = 430.53*103j/mol

this energy is for 1 mol so for 1 molecule

E = 430.5*103/Na                         (Na is avagadro number)

E = 7.2*10-19joule

 

energy of radiation = hc/lamda = 7.8*10-19joules

here h is planks constant , c is speed of light , lamda is wavelength)

 

energy given through radiation is more than dissotion so excess  energy

will come in form of KE of dissociated atoms ...

 

dE = axcess energy = (7.8-7.2)*10-19 = 0.6*10-19

 

(dE/E)*100 = 8%

so 7.8% of energy is converted to KE...